Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

sports league

Expert replies
by CSASHISHPANDAY » Wed Jul 25, 2012 9:38 am
A certain sports league has ten teams in its Western Conference and eight teams in its Eastern Conference. At the end of the season, two teams, one from each conference, play in The Big Game. Two teams in each conference are located in Texas. If each team in each conference has an equal probability of making it to The Big Game, and if each team in The Big Game has an equal probability of winning that game, what is the probability that a team from Texas will win The Big Game?

(A)3/20


(B)2/9


(C) 9/40


(D)27/80


(E) 5/16

OA and source after discussion
Join the discussion
Source: — Problem Solving |

by CSASHISHPANDAY » Wed Jul 25, 2012 9:54 am
here is an answer of the question https://www.beatthegmat.com/probability- ... 86396.html

but i m unable to understand how the probability of Texas team reaches final from West,p(w)is 2/10.
Join the discussion

by niketdoshi123 » Wed Jul 25, 2012 10:19 am
CSASHISHPANDAY wrote: but i m unable to understand how the probability of Texas team reaches final from West,p(w)is 2/10.
Two teams in each conference are located in Texas
So 2 out of 10 teams in Western Conference are located in Texas

Also, each team in each conference has equal probability of making it to The Big Game.
So probability that a team in western conference makes it to the final = 1/10

Since there are two teams from Texas, required probability = 2*1/10
Join the discussion

by GMATGuruNY » Wed Jul 25, 2012 10:28 am
CSASHISHPANDAY wrote:A certain sports league has ten teams in its Western Conference and eight teams in its Eastern Conference. At the end of the season, two teams, one from each conference, play in The Big Game. Two teams in each conference are located in Texas. If each team in each conference has an equal probability of making it to The Big Game, and if each team in The Big Game has an equal probability of winning that game, what is the probability that a team from Texas will win The Big Game?

(A)3/20


(B)2/9


(C) 9/40


(D)27/80


(E) 5/16

OA and source after discussion
Case 1: A Western team from Texas wins
P(a Western team from Texas is in the Big Game) = 2/10. (Of the 10 Western teams, 2 are from Texas.)
P(this team wins) = 1/2. (Since each team in the Big Game has an equal chance of winning.)
Since we want both of these events to happen, we multiply the fractions:
2/10 * 1/2 = 1/10.

Case 2: An Eastern team from Texas wins
P(an Eastern team from Texas is in the Big Game) = 2/8. (Of the 8 Eastern teams, 2 are from Texas.)
P(this team wins) = 1/2. (Since each team in the Big Game has the same chance of winning.)
Since we want both of these events to happen, we multiply the fractions:
2/8 * 1/2 = 1/8

Since either case will yield a favorable outcome,we ADD the probabilities:
1/10 + 1/8 = 4/40 + 5/40 = 9/40.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion